# What is ‘Typological Diversity’ in NLP?

### v1

@@ 5.2 Inclusion of Typological Features
Because lang2vec distance calculations are based on feature vectors, our previous analysis does not provide information regarding which typological feature values are actually included in the language set. For example, it does not tell us which word order variations are covered in a given language selection. This can be useful, because covering more typological feature values in an evaluation set means that robustness towards previously unseen typological characteristics is increased. Firstly, we look into the inclusion of individual typological feature values per language selection from each paper. We use the Grambank (Skirgård et al., 2023) database for this, because is has high coverage for grammatical features and is currently actively maintained. For each feature in Grambank, we count whether all possible, non-missing feature values are represented in the paper’s language selection. [^4] Note that for our purposes, we treat ? as a missing value. That is, for a feature such as GB020: Are there definite or specific articles?, we count whether all non-missing options are represented by the particular set of languages from a paper and divide this by the total number of features. The spread of these values is shown in the bottom right graph in Figure 5. We treat missing languages the same as in the MPSD calculations; we discard them during calculation, but we do not discard the entire paper. Twelve papers are affected by this. The average Grambank feature value inclusion is 0.73.
While the accumulation of absolute feature values covered in a language selection provides useful robustness insights, it does not capture that languages consist of combinations of features that are not entirely independent. Skirgård et al. (2023) show how language families can reliably be visualized as distinct groupings in Grambank’s typological design space using principle component analysis (PCA). Similarly, in Figure 6 we show the PCA plot for the paper that has the highest Grambank feature inclusion of all papers: Gutierrez-Vasques et al. (2021). While their language sample covers much of the Grambank’s typological space, we observe a skew towards certain languages (bottom right), and a lack of representation for others (e.g., bottom left). This means that even the paper in our survey with the highest feature value inclusion contains challenges when it comes to complete and fair evaluation.
@@ 5.3 Number of Languages
As shown in Figure 3, the number of included languages varies across papers. Should one aim to evaluate on as many languages as possible? What is the effect of selecting many (similar) languages? Figure 7 and 7 show the relationship between the number of languages and the MPSDs, and the absolute typological feature value inclusion, respectively. We find that smaller language sets (0-10 languages) exhibit considerable variation in terms of average syntactic distances, while the larger language sets (30+ languages) seem consistent between 0.6 and 0.7. This implies that adding more languages does not necessarily raise the average syntactic distance much. In NLP, adding more languages typically means adding more similar languages, since it is easier to incorporate existing datasets than it is to create new ones. As a result, the average performance will also be skewed towards these similar languages in evaluation (see Section 6). Furthermore, we observe that up to a certain point, including more languages implies that more typological feature values are covered. However, this increase flattens at approximately 40 languages. Importantly, this highlights the fact that simply adding more languages to an experimental study is not by itself a contribution in terms of typological generalizability – one must take care which languages are included. To illustrate this, we highlight the MKQA dataset Longpre et al. (2021). This dataset includes 25 languages, more than twice the median, but has a rather average MPSD of 0.61 (and a Grambank feature inclusion of 0.89). This is in part due to its inclusion of several typologically similar languages: {English, German, Dutch}, {Cantonese, Mandarin}, {Spanish, Portuguese} and {Swedish, Danish, Norwegian}.
@@
Table 2: XTREME-R results grouped by inflection type. Overall refers to the average over all languages. By Feature is the average of the WALS feature averages, excluding languages for which there is no coverage. The delta shows the difference of Overall and By F. Tasks that have coverage for all their included languages are on top, those partially covered are in the bottom portion. Morphological inflection types are: Aff = affixing, Suf = suffixing, Pre = prefixing and NA referring to Not Available (in WALS). The highest and the lowest number of languages per grouping are highlighted. The (italicized) number refers to the number of languages in a particular subset. Metrics: ♠ = Accuracy, ★ = mAP@20, ♠ = F1.
@@ 6 Downstream Implications
Inconsistencies in ‘typologically diverse’ language selection can have an effect on downstream evaluation. Anastasopoulos (2019) demonstrates that, when a language set contains a skew towards a certain group of languages (in their case: language family), simply taking the average over all languages for evaluation can give an overestimation of performance. Naturally, reporting all metrics for all languages separately is preferable, but often not practical due to space constraints: these results, if included, are generally placed in an appendix.
Beyond language family, we provide additional insights into the effects of skewed language selection for model evaluation. We do this by averaging based on typological groupings. Specifically, we look at XTREME-R Ruder et al. (2021), a popular [^5] The original XTREME dataset has almost 800 citations as of writing, XTREME-R, which is an extension of the original, more than 100. ‘typologically diverse’ multilingual benchmark covering several tasks and domains. XTREME-R contains and expands upon ten existing datasets. As a result, not all tasks are available in all languages; the number of languages per task ranges from 7 to 48, while across tasks there are 50 unique languages in total. [^6] See Appendix E, Figure 9. The authors provide macro averages by language family and by task.
We group these languages by the ‘Prefixing vs. Suffixing in Inflectional Morphology’ and ‘Order of Subject, Object and Verb’ (word order) features, as provided by WALS (Dryer, 2013). These features have high language coverage, which will make our analysis as accurate as possible. Table 2 shows groupings and coverage of the inflectional feature, the word order analysis is included in Appendix D, Table 5. ‘Strongly Suffixing’ is the most represented group for all tasks by quite a margin (the orange column), with ‘Little Affixing’ coming in second, being covered in all but one task. Interestingly, the performance differences between ‘Strongly Suffixing’ and the others are not consistently higher or lower, despite the big difference in the number of languages that exhibit these phenomena. ‘Strongly Prefixing’ is not explicitly covered in any task. The delta average column shows the difference between the macro average over all languages and over all features. Here, we see similar patterns as Anastasopoulos (2019) has shown; skewed performance numbers ranging from substantially worse (Mewsli-X), to even slightly better, depending on the model (MLQA). These patterns were not uncovered by averaging based on language family or over all languages.
@@ 7 Insights and Recommendations
We systematically analyzed claims of ‘typological diversity’ in NLP research. We approximated diversity in terms of average syntactic language distance (MPSD) and absolute typological feature value inclusion in Grambank. Our analysis shows that (1) there is no consistent definition or methodology when making ‘typological diversity’ claims, (2) our estimates of typological diversity exhibit a considerable variation across papers and (3) aggregated results can give distorted views of multilingual performance estimates.
We recommend future approaches to include an operationalization of ‘typological diversity’ when making such claims. Additionally, we recommend adding an empirical justification, such as the feature value coverage or PCA plots we have shown. This has the potential to benefit multilingual NLP, as it enables more fine-grained insights into typological challenges of multilingual language modelling.
@@ Appendix A Paper Retrieval Details
Appendix B Long Tail of Included Languages
@@
Appendix C ‘Typologically Diverse’ Datasets
Paper Dataset name ||||L|normal-||| (normal-$\downarrow$) MPSD GB Task(s) or Topic
@@
Table 4: Papers that have a claim and introduce a dataset. L refers to the number of languages regarding the claim, with † indicating that one language does not have an ISO-639-3 code. MPSD is the Mean Pairwise Distance of syntactic URIEL features per language pair, with § indicating that at least one language has low ($<5\%$) syntactic feature coverage. GB refers to the feature coverage (i.e., are all possible, non-missing values for a feature covered) of the included languages in Grambank. Here, ‡ refers to at least one language not being covered in Grambank, these languages are excluded from the coverage metric. The highest and lowest values for MPSD and GB are highlighted. Abbreviations: QA = Question Answering, RE = Reasoning, NLI = Natural Language Inference, IR = Information Retrieval, ASR = Automatic Speech Recognition, MT = Machine Translation.
Appendix D XTREME-R Results by Word Order
@@
Table 5: XTREME-R task and model results, grouped by word order. Overall refers to the average over all languages. By Feature is the average of the WALS feature averages, excluding languages for which there is no coverage. The delta shows the difference of Overall and By F. Tasks that have coverage for all their included languages are on top, those partially covered are in the bottom portion. The word orders are listed in the following columns, with NDO being No Dominant Order and NA referring to Not Available (in WALS). The highest and the lowest number of languages per grouping are highlighted. The (italicized) number refers to the number of languages in a particular subset. Metrics: ♠ = Accuracy, ★ = mAP@20, ♠ = F1.
Appendix E XTREME-R Grambank Feature Inclusion
@@
Figure 9: PCA Grambank feature inclusion per XTREME-R subtask. Where applicable, we cite the original dataset that XTREME-R includes or used as a starting point to expand upon.

### v3

@@ 5.2 Inclusion of Typological Features
Because lang2vec distance calculations are based on feature vectors, our previous analysis does not provide information regarding which typological feature values are actually included in the language set. For example, it does not tell us which word order variations are covered in a given language selection. This information can be useful though, because covering more typological feature values in an evaluation set means that robustness towards previously unseen typological characteristics is increased. Following the saturation measure by Miestamo et al. (2016), we look into the inclusion of individual typological feature values per language selection from each paper. We use the Grambank [^9] https://grambank.clld.org/ (Skirgård et al., 2023b) database, because it has high coverage for grammatical features and is currently actively maintained. For each feature in Grambank, we count whether all possible, non-missing feature values are represented in the paper’s language selection. [^10] Note that for our purposes, we treat ? as a missing value. That is, for a feature such as GB020: Are there definite or specific articles?, we count whether all non-missing options are represented by the particular set of languages from a paper and divide this by the total number of features. The spread of these values is shown in the bottom right graph in Figure 5. We treat missing languages the same as in the MPSD calculations; we discard them during calculation, but we do not discard the entire paper. Twelve papers are affected by this. The average Grambank feature value inclusion is 0.73.
While the accumulation of absolute feature values covered in a language selection provides useful robustness insights, it does not capture that languages consist of combinations of features that are not entirely independent. Skirgård et al. (2023a) show how language families can reliably be visualized as distinct groupings in Grambank’s typological design space using principle component analysis (PCA). Similarly, in Figure 6 we show the PCA plot for the paper that has the highest Grambank feature inclusion of all papers: Gutierrez-Vasques et al. (2021). While their language sample covers much of the Grambank’s typological space, we observe a skew towards certain languages (bottom right), and a lack of representation for others (e.g., bottom left). This means that even the paper with the highest feature value inclusion in our survey contains challenges when it comes to complete and fair evaluation.
@@ 5.3 Number of Languages
As shown in Figure 3, the number of included languages varies across papers. Should one aim to evaluate on as many languages as possible? What is the effect of selecting many (similar) languages? Figure 7 and 7 show the relationship between the number of languages and the MPSDs, and the absolute typological feature value inclusion, respectively. We find that smaller language sets (0-10 languages) exhibit considerable variation in terms of average syntactic distances, while the larger language sets (30+ languages) seem consistent between 0.6 and 0.7. This implies that adding more languages does not necessarily raise the average syntactic distance much. In NLP, adding more languages typically means adding more similar languages, since it is easier to incorporate existing datasets than it is to create new ones. As a result, the average performance will also be skewed towards these similar languages in evaluation (see §6). Furthermore, we observe that up to a certain point, including more languages implies that more typological feature values are covered. However, this increase flattens at approximately 40 languages. Importantly, this highlights the fact that simply adding more languages to an experimental study is not by itself a contribution in terms of typological generalizability – one must take care which languages are included. To illustrate this, we highlight the MKQA dataset Longpre et al. (2021). This dataset includes 25 languages, more than twice the median, but has a rather average MPSD of 0.61 (and a Grambank feature inclusion of 0.89). This is in part due to its inclusion of several typologically similar languages: {English, German, Dutch}, {Cantonese, Mandarin}, {Spanish, Portuguese} and {Swedish, Danish, Norwegian}.
@@
Table 2: XTREME-R results grouped by inflection type (WALS 26A). Overall refers to the average over all languages. By Feature is the average of the WALS feature averages, excluding languages for which there is no coverage. The delta shows the difference of Overall and By F. Tasks that have coverage for all their included languages are on top, those partially covered are in the bottom portion. Morphological inflection types are: Aff = affixing, Suf = suffixing, Pre = prefixing and NA referring to Not Available (in WALS). The highest and the lowest non-zero number of languages per grouping are highlighted. The (italicized) number refers to the number of languages in a particular subset. Metrics: ♠ = Accuracy, ★ = mAP@20, ♠ = F1.
@@ 6 Downstream Implications
Inconsistencies in ‘typologically diverse’ language selection can have an effect on downstream evaluation. While reporting all metrics for all languages separately is preferable, space constraints often lead to reporting averages across languages in multilingual evaluation. However, Anastasopoulos (2019) demonstrates that when a language set contains a skew towards a certain language family, simply taking the micro average over all languages for evaluation gives an overestimation of performance. While family can serve as a proxy for typological diversity, we here present an analysis based directly on typological characteristics. In this way, we gain more fine-grained insights into the effects of typologically skewed language selection for model evaluation.
Specifically, we use XTREME-R Ruder et al. (2021), a popular [^11] The original XTREME dataset has almost 800 citations as of writing, XTREME-R, which is an extension of the original, more than 100. ‘typologically diverse’ multilingual benchmark covering several tasks and domains. XTREME-R contains and expands upon ten existing datasets. As a result, not all tasks are available in all languages; while there are 50 unique languages in total, the number of languages per task ranges from 7 to 48.121212See Appendix F, Figure 9. We group these languages by the ‘Prefixing vs. Suffixing in Inflectional Morphology’ (26A) and ‘Order of Subject, Object and Verb’ (81A) features, as provided by WALS (Dryer, 2013b, a). These features have high language coverage, which makes our analysis as comprehensive as possible. Table 2 shows groupings and coverage of the inflectional feature (26A). The word order analysis is included in Appendix E, Table 5.
Firstly, we see that XTREME-R contains a skew in terms of morphological inflection. For all tasks, the majority of languages are strongly suffixing (see orange cells). Some feature values, such as equal prefixing and suffixing and weak prefixing are underrepresented, with at most one language per subtask. Remarkably, strong prefixing does not appear in XTREME-R at all. This implies that one should be careful with the implication of generalizability that evaluating on a ‘typologically diverse’ dataset gives.
The delta column shows the difference between the macro average over all languages and the macro average per feature value. Here, we observe similar patterns as in Anastasopoulos (2019), namely that estimations of multilingual model performance vary considerably when accounting for typological imbalances. We suggest that drawing and implying generalizable conclusions about typologically diverse evaluation should ideally be supported by showing that performance holds across a range of typological properties.
@@ 7 Insights and Recommendations
We systematically analyze claims of ‘typological diversity’ in NLP research. We approximate diversity in terms of average syntactic language distance (MPSD) and absolute typological feature value inclusion in Grambank. Our analysis shows that (1) there is no consistent definition or methodology when making ‘typological diversity’ claims, (2) estimates of typological diversity exhibit a considerable variation across papers and (3) aggregated results can give distorted views of multilingual performance estimates.
We recommend future approaches to include an operationalization of ‘typological diversity’ when making such claims. This can be in terms of generalizability or regarding a particular typological phenomenon of interest. Additionally, we recommend adding an empirical justification, especially when claims relate to generalizability. The feature value inclusion, MPSD or PCA plots we have shown are examples of these justifications. Including these has the potential to benefit multilingual NLP, as it enables more fine-grained insights into typological challenges of language modelling.
@@
Another limitation is that our search is based only on the title and abstract of papers. There are almost certainly papers that contain a claim in other sections. However, we chose this route since we wanted to (1) cast a wide net, (2) perform some pre-filtering of papers, assuming that a mention of ‘typological diversity’ in the title or abstract is a good indicator that it is a prominent aspect of that paper, and (3) make sure that the data we could get was of good quality, which is a lot harder with (old) pdf files for example.
Our arguments for assessing linguistic typology directly in NLP, rather than phylogeny and geography as proxies, relates to claims of ‘typological diversity’. We acknowledge that geographical and genealogical information could be useful for other purposes.
@@ Appendix A Paper Retrieval Details
Appendix B Diversity Metrics
Mean Pairwise Distance (MPD)
For a language sample $L$ and feature vectors $V$, we retrieve $v_{l}$ for $l\in L$. We then calculate the mean, pairwise Euclidean distances between all pairs in $L$:
$MPD(L)=\frac{1}{|L|(|L|-1)}\sum_{i,j\in L,i\neq j}\lVert v_{i}-v_{j}\rVert_{2}$
where the feature vectors from $V$ consists of the geography, genetic or syntactic vectors in lang2vec (as shown per paper in Figure 5).
Feature Value Inclusion (FVI)
[…]
$FVI(L)=\frac{1}{|\mathcal{F}|}\sum_{i=1}^{|\mathcal{F}|}\frac{|\{v_{l}[F_{i}]:l\in L\}|}{|\mathcal{V}_{F_{i}}|}$
Appendix C Long Tail of Included Languages
@@
Appendix D ‘Typologically Diverse’ Datasets
Paper Dataset name ||||L|||| ($\downarrow$) MPSD FVI Task(s) or Topic
@@
Table 4: Papers that have a claim and introduce a dataset. L refers to the number of languages regarding the claim, with † indicating that one language does not have an ISO-639-3 code. MPSD is the Mean Pairwise Distance of syntactic URIEL features per language pair, with § indicating that at least one language has low ($<5\%$) syntactic feature coverage. FVI refers to the Grambank feature value inclusion of the languages in the dataset: are all possible, non-missing values for a feature in Grambank included using the language selection. Here, ‡ refers to at least one language not being covered in Grambank, these languages are excluded from the coverage metric. The highest and lowest values for MPSD and FVI are highlighted. Abbreviations: QA = Question Answering, RE = Reasoning, NLI = Natural Language Inference, IR = Information Retrieval, ASR = Automatic Speech Recognition, MT = Machine Translation.
Appendix E XTREME-R Results by Word Order
@@
Table 5: XTREME-R task and model results, grouped by word order (WALS 81A). Overall refers to the average over all languages. By Feature is the average of the WALS feature averages, excluding languages for which there is no coverage. The delta shows the difference of Overall and By F. Tasks that have coverage for all their included languages are on top, those partially covered are in the bottom portion. The word orders are listed in the following columns, with NDO being No Dominant Order and NA referring to Not Available (in WALS). The highest and the lowest number of languages per grouping are highlighted. The (italicized) number refers to the number of languages in a particular subset. Metrics: ♠ = Accuracy, ★ = mAP@20, ♠ = F1.
Appendix F XTREME-R Grambank Feature Value Inclusion
@@
Figure 9: PCA Grambank feature value inclusion per XTREME-R subtask. Where applicable, we cite the original dataset that XTREME-R includes or used as a starting point to expand upon.